if we hav to brought a mass from infinitey to the vertices of equaterial triangle minimun work done by ext agent is????
we hav a mass of 1kg at each vertices and each side is 0.1m long

Grade:12

1 Answers

Askiitians Expert Bharath-IITD
14 years ago

Dear Jauneet,

The Work done for bringing the masses from infinity to the vertices of the euilateral triangle is nothing but sum of potential energies of

1. The first mass which is zero because there is no other mass.

2. Second mass is the potential energy between the first mass and the present one

U12= -G m2/r where r is the of the equilateral triangle.

3. Third mass (i) withrespect to the first mass U13 = -G m2/r

                    (ii) withrepect to the second mass U23 = -G m2/r

Thus the total work done is given by W = UT = U12 + U13 + U23 = - 3* G m2/r = 20.022 * 10-9

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best !!!

 


Regards,

Askiitians Experts

Adapa Bharath